What is the solution for y=117-x and y=24.5-x? Thank you
Answer by stanbon(75887)

What is the solution for
y=117-x
and
y=24.5-x
-----
Note: There is no x value that can make both of those true.
----
Showing the algebra:
Substitute y for y to get:
117-x = 24.5-x
----
Add x to both sides to get:
117 = 24.5
-----
Obviously wrong.
Why??? Because we assumed those two "x's" were the same.
They cannot be.
====================
Answer: No solution
====================
Cheers,
Stan H.